# Efficient Beam Selection for ISAC in Cell-Free Massive MIMO via Digital Twin-Assisted Deep Reinforcement Learning

### v1

@@ VI Conclusion
This paper investigated the problem of joint target detection and proposed a novel beam selection method for cell-free ISAC systems. We employed offline DRL based on a dueling DDQN network and meticulously designed reward shaping function and sinusoidal embedding to enable effective beam selection while ensuring adherence to predetermined false alarm rates and target detection threshold. To improve the learning efficiency of the offline DRL agent, we proposed a DT-assisted framework based on cGAN to enrich the data diversity and incorporated the CQL function to address the overestimation problem. Numerical results demonstrated that our proposed method offers significant advantages in the beam selection task, reducing the number of interactions by 80% while maintaining performance comparable to online DRL. Furthermore, our proposed method exhibits robustness across various transmitting power levels, false alarm rate probabilities, and target velocities.

### v2

@@ VI Conclusion
This paper investigated the problem of joint target detection and proposed a novel beam selection method for cell-free ISAC systems. We employed offline DRL based on a dueling DDQN network and meticulously designed reward shaping function and sinusoidal embedding to enable effective beam selection while ensuring adherence to predetermined false alarm rates and target detection threshold. To improve the learning efficiency of the offline DRL agent, we proposed a DT-assisted framework based on cGAN to enrich the data diversity and incorporated the CQL function to address the overestimation problem. Numerical results demonstrate that our proposed method offers significant advantages in the beam selection task, reducing the number of interactions by 80% while maintaining performance comparable to online DRL. Furthermore, our proposed method exhibits robustness across various transmitting power levels, false alarm rate probabilities, and target velocities.
In future works, we plan to exploit communication signals to assist target detection and enable joint ISAC optimization. Multi-agent learning frameworks will be investigated to address the exponential growth of the action space as the number of users increases. Moreover, integrating DT with multi-agent DRL represents another promising direction for enhancing system scalability and coordination efficiency.